OpenStudy (anonymous):

Multiply: (8-3i)(5-4i)

OpenStudy (anonymous):

http://www.twiddla.com/solved

OpenStudy (anonymous):

38-47i

OpenStudy (anonymous):

28-47i

OpenStudy (anonymous):

soory..28-47i

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Foid 40-32i-15i+12i^2 i^2=-1 40-47i+(12*-1)=40-47i -12 = 28 -47i

OpenStudy (anonymous):

You can "foil" or you can use traditional top down multiplication methods, which are much easier in my opinion, lol

OpenStudy (anonymous):

only foid

OpenStudy (anonymous):

What is foid....lol

OpenStudy (anonymous):

time

OpenStudy (anonymous):

I thought it was foil, lol

OpenStudy (anonymous):

(8*5)-(8*-4i)-(3i*5)+(-3i*-4i)

OpenStudy (anonymous):

That's FOIL....first, outer, inner, last...but, there's always more than one way of course....

OpenStudy (anonymous):

FOIL is popular, but it isn't the traditional form of multiplication method....It's more like a "way to remember how to multiply in case you forget" method, lol